OpenStudy (anonymous):

f(x)=-4x-2 find f(2)

OpenStudy (anonymous):

just replace x with 2 so f(2) = -4*2-2 = -8-2 = -10

OpenStudy (anonymous):

thank you appreciate for that .

OpenStudy (anonymous):

just subtitute x with 2 :)